Opened 10 years ago

Last modified 10 years ago

#423 new flight processing

BB11_01, flight day 272c/2011, Brecon Beacons

Reported by: adbe Owned by: jaho
Priority: alpha 4 high Milestone: 2011 data processing completion
Component: Processing: general Keywords:
Cc: Other processors: adbe,knpa

Description (last modified by adbe)

Data location: ~arsf/arsf_data/2011/flight_data/uk/BB11_01-2011_272c_Brecon_Beacons

Data arrived from ARSF via network transfer on 06/10/2011

BB11_01 Scientific objective:

BB11_01 Priority: a4h

PI: Shaun Lewis

Sensors:

Camera (23/11/2011)
Eagle (23/11/2011)
Hawk (23/11/2011)
LiDAR (23/11/2011)

Change History (22)

comment:1 Changed 10 years ago by adbe

Second part of 272a.

comment:2 Changed 10 years ago by adbe

Processed navigation

comment:3 Changed 10 years ago by adbe

Started hyperspectral processing

comment:4 Changed 10 years ago by jaho

Started LiDAR processing.

comment:5 Changed 10 years ago by jaho

LiDAR ready for delivery check.

comment:6 Changed 10 years ago by adbe

Hyperspectral ready for delivery check

comment:7 Changed 10 years ago by jaho

Started Camera processing.

comment:8 Changed 10 years ago by jaho

Removed image 107. Should be in 272a.

comment:9 Changed 10 years ago by jaho

Camera ready for delivery check.

comment:10 Changed 10 years ago by emca

  • Description modified (diff)
  • Summary changed from BB11_01, flight day 272a/2011, Brecon Beacons to BB11_01, flight day 272c/2011, Brecon Beacons

Delivery checked LiDAR data.

bng DEM is not in WGS84, so this should be removed from the name (& header and readme). Same applies to the DEM in 272a.

Otherwise ready to deliver.

comment:11 Changed 10 years ago by jaho

Fixed DEMs' names. Ready to deliver.

comment:12 Changed 10 years ago by knpa

Delivery checked RCD.

Perfect! Ready to deliver.

comment:13 Changed 10 years ago by knpa

Gary says that the PI should be able to get their own vectors via PGA.
Do not wait for vectors to make delivery but deliver with caveat that we have not made this usual check.

comment:14 Changed 10 years ago by emca

Hyperspectral delivery check complete.

One small change required to readme - bottom of page 8 the text is going off the side of the page.

Otherwise data is all good, ready to deliver.

comment:15 Changed 10 years ago by adbe

Made change to hyperspectral ReadMe

comment:16 Changed 10 years ago by adbe

Camera, Eagle, Hawk and LiDAR on disk 117 and ready to deliver on 08/11/2011 pending payment

comment:17 Changed 10 years ago by jaho

Reprocessed LiDAR with pitch offset -0.00299419. Ready for delivery check under BB11_01-272c-lidar-20111123/.

comment:18 Changed 10 years ago by adbe

Lidar Delivery Check:

No problems, ready to deliver

comment:19 Changed 10 years ago by jaho

Updated LiDAR ReadMe to mention vectors.

comment:20 Changed 10 years ago by adbe

  • Description modified (diff)

Camera, Hyperspectral and LiDAR delivered to Shaun Lewis on 23/11/2011 on disk 117

comment:21 Changed 10 years ago by mark1

Hyperspectral hawk bad pixel files and XML files have been re-created and re-delivered to PI (via email). This corrects sample number in bad pixel text files and Eagle IFOV value in XML and extra comments included.

comment:22 Changed 10 years ago by adbe

Hyperspectral sct's

lineeaglehawk
8-0.01-0.02
9-0.08-0.08
10-0.05-0.04
11-0.04-0.04
120.00-0.02
Note: See TracTickets for help on using tickets.